# Leveraging Priors via Diffusion Bridge for Time Series Generation

### v1

@@ Abstract
Time series generation is widely used in real-world applications such as simulation, data augmentation, and hypothesis test techniques. Recently, diffusion models have emerged as the de facto approach for time series generation, emphasizing diverse synthesis scenarios based on historical or correlated time series data streams. Since time series have unique characteristics, such as fixed time order and data scaling, standard Gaussian prior might be ill-suited for general time series generation. In this paper, we exploit the usage of diverse prior distributions for synthesis. Then, we propose TimeBridge, a framework that enables flexible synthesis by leveraging diffusion bridges to learn the transport between chosen prior and data distributions. Our model covers a wide range of scenarios in time series diffusion models, which leverages (i) data- and time-dependent priors for unconditional synthesis, and (ii) data-scale preserving synthesis with a constraint as a prior for conditional generation. Experimentally, our model achieves state-of-the-art performance in both unconditional and conditional time series generation tasks.
Introduction
Synthetic time series have expanded their applications by serving roles in data augmentation (Wen et al. 2021), privacy-preserving synthesis (Wang et al. 2024), continuous data compression (Zhou et al. 2023), and simulation of specific hypothesis (Xia et al. 2024) for the data streams. With the advancement of deep generative architectures, time series generative models have been actively investigated such as variational autoencoders (VAE) (Desai et al. 2021; Naiman et al. 2024), generative adversarial networks (GANs) (Yoon, Jarrett, and Van der Schaar 2019; Jeon et al. 2022; Xia et al. 2024), and diffusion models (Ho, Jain, and Abbeel 2020; Karras et al. 2022).
Diffusion models are well-known for producing high-quality and diverse time series data samples (Tashiro et al. 2021; Alcaraz and Strodthoff 2023). The effectiveness of these models relies on iteratively estimating small steps of the reverse trajectory, which corresponds to the forward process towards standard Gaussian priors (Song et al. 2021).
@@ Introduction
However, modeling diffusion with constraints is not flexible due to the predetermined mapping from noise to data (Harvey et al. 2022; Du et al. 2023). In the time series domain, Coletta et al. (2024) highlighted that generating data under specific bull or bear market patterns requires tailored embedding or penalty functions for sampling guidance. Additionally, the sequential order and data scales (Kim et al. 2022) must be considered when modeling the statistics of historical data and time dependencies (Zeng et al. 2023).
@@
Figure 1: Illustration of the proposed time series diffusion bridge framework focusing on prior design for diverse tasks.
Recently, altering standard Gaussian noise in diffusion models has been explored by modeling the data-dependent distribution in both forward and reverse processes (Han, Zheng, and Zhou 2022; Yu et al. 2024). Specifically, in the time series domain, the noising process should consider not only the diffusion time steps but also the time stamps of temporal data to maintain the continuity of generated data and time dependency (Biloš et al. 2023; Chen et al. 2023a). In this paper, we adopt the diffusion bridge (Schrödinger 1932; De Bortoli et al. 2021; Zhou et al. 2024) to utilize diverse diffusion priors. By learning the optimal transport between the data distribution and fixed priors, diffusion bridges make use of fixed priors and have proven effective in conditional tasks such as image-to-image translation (Li et al. 2023) and text-to-speech synthesis (Chen et al. 2023b).
To push further, we investigate the various prior designs proper for time series tasks. Then, we propose TimeBridge, a flexible time series generation framework that addresses multiple time series synthesis with the prior design of diffusion bridge, as shown in Figure 1. We then deeply investigate various experimental settings from unconditional to conditional tasks, achieving state-of-the-art performance across multiple datasets, as illustrated in Figure 2. Our contributions to various situations are summarized as follows:
  • We explore the prior selections on time series and propose TimeBridge, a diffusion bridge model specifically designed to leverage the priors for time series synthesis.
  • For unconditional generation, we propose prior distributions tailored to time series synthesis by leveraging data- and time-dependent properties.
@@
  • For conditional generation, we integrate the constraints into the pair-wise prior to maintaining data scales. We also introduce point-preserving sampling to preserve constraints, extending its application to imputation tasks.
@@
Figure 2: Performance measures (lower is better) on different tasks and sampling time (lower is better) of time series diffusion models. The proposed TimeBridge achieves both generalization in various tasks and sampling efficiency.
Preliminaries
@@
Consider the multivariate time series ${\mathbf{x}}$ defined within the sample space $\mathcal{X}=\mathbb{R}^{d\times\tau}$, where $d$ represents the dimensionality of the variables and $\tau$ denotes the time length. For each index $i\in\{1,\ldots,N\}$ among the $N$ samples, each time series is denoted as ${\mathbf{x}}^{i}=({\bm{x}}^{i}(1),\ldots,{\bm{x}}^{i}(\tau))\in\mathcal{X}$, given the data ${\bm{x}}^{i}(k)\in\mathbb{R}^{d}$ for each time step $k\in\{1,\ldots,\tau\}$. For diffusion steps $t\in\{0,\ldots,T\}$, we denote intermediate time series samples as ${\mathbf{x}}_{t}$, and for a specific instance $i$ as ${\mathbf{x}}_{t}^{i}$.
Diffusion process.
@@
In standard diffusion models (Ho, Jain, and Abbeel 2020), the diffusion process is constructed by gradually injecting noise into samples ${\mathbf{x}}_{0}$ drawn from the data distribution $p_{0}$, forwarding them into a standard Gaussian distribution $p_{T}=\mathcal{N}(\bm{0},\bm{I})$. The corresponding stochastic differential equation (SDE) (Song et al. 2021) is:
$d{\mathbf{x}}_{t}={\mathbf{f}}({\mathbf{x}}_{t},t)dt+g(t)d{\mathbf{w}}_{t},$ (1)
@@
where ${\mathbf{f}}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ represents the drift function, $g:[0,T]\rightarrow\mathbb{R}$ denotes the diffusion coefficient, and ${\mathbf{w}}_{t}$ is the Wiener process. The reverse SDE is formulated as follows:
$d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},$ (2)
@@
where $p({\mathbf{x}}_{t})$ refers to the probability density of ${\mathbf{x}}_{t}$ (Anderson 1982) and $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})$ is a score function to match.
Diffusion bridge.
@@
The aforementioned standard diffusion models cannot be extended to more general tasks due to their inflexibility in choosing a prior distribution. Instead, to construct the diffusion process towards any endpoint, Doob’s h-transform (Doob and Doob 1984) can be applied to Equation (1). With a fixed endpoint ${\mathbf{y}}$, the forward process is:
$\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g(t)^{2}{\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)]dt+g(t)d{\mathbf{w}}_{t},\\ {\mathbf{x}}_{0}\sim q_{data}({\mathbf{x}}),{\mathbf{x}}_{T}={\mathbf{y}},\end{split}$ (3)
@@
where ${\mathbf{h}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}$. For a given data distribution $q_{data}({\mathbf{x}},{\mathbf{y}})$, the diffusion bridge (Schrödinger 1932; De Bortoli et al. 2021) aims to transport ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{T}$ where $({\mathbf{x}}_{0},{\mathbf{x}}_{T})=({\mathbf{x}},{\mathbf{y}})\sim q_{data}({\mathbf{x}},{\mathbf{y}})$ while following above forward process. Similar to the standard diffusion model, we can construct the reverse SDE as follows:
$\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}({\mathbf{s}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)-\\ {\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T))]dt+g(t)d\bar{{\mathbf{w}}}_{t},\end{split}$ (4)
@@
${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}.$ (5)
For training, it is sufficient to learn the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$.
The Schrödinger bridge (Schrödinger 1932) is a closely related concept, focusing on the path optimizations between two arbitrary distributions. Diffusion bridge models have become popular due to their abilities (i) to flexibly map in various domains, including image-to-image translation (Li et al. 2023; Shi et al. 2022) and text-to-speech synthesis (Popov et al. 2021; Chen et al. 2023b), and (ii) to achieve efficient sampling with fewer diffusion steps (Zhou et al. 2024; Chen et al. 2023b).
Reparametrization of diffusion bridge.
@@
To approximate the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$ in Equation (5), Zhou et al. (2024) recently proposed a Denoising Diffusion Bridge Model (DDBM), integrating the diffusion bridge model in the framework of standard diffusion models. DDBM enables tractable marginal sampling of ${\mathbf{x}}_{t}$, by designing ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are the noise schedule functions and ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$. For the variance-preserving (VP) schedule, the marginal distribution at time $t$ is as follows:
$q({\mathbf{x}}_{t}|{\mathbf{x}}_{0},{\mathbf{x}}_{T})=\mathcal{N}\left(\hat{\bm{\mu}}_{t},\hat{\sigma}^{2}_{t}\mathbf{I}\right),$ (6)
$\hat{\bm{\mu}}_{t}=\textstyle\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}{\mathbf{x}}_{0}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$ (7)
$\hat{\sigma}^{2}_{t}=\textstyle\sigma^{2}_{t}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$ (8)
@@
where signal-to-ratio (SNR) is defined as $SNR_{t}={\alpha^{2}_{t}}/{\sigma^{2}_{t}}$. Since the mean $\hat{\bm{\mu}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}$ at time $t$, the data scale is preserved for translation. With the reparametrization of Elucidating Diffusion Models (EDM) (Karras et al. 2022), we can match the score as follows:
$\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})\approx{\mathbf{s}}({D_{\theta}},{\mathbf{x}}_{t},t,{\mathbf{x}}_{T},T)$ (9)
@@
Problem Statement
In this section, we cover two types of time series synthesis tasks, i.e., unconditional and conditional generations.
Unconditional generation considers the total dataset containing $N$ samples, denoted as $\mathcal{D}=\{{{\mathbf{x}}^{i}}\}_{i=1}^{N}$. Our goal is to train a deep generative model to mimic the input dataset and yield synthetic samples statistically similar to a given time series dataset. We aim to make the distribution $p(\hat{\mathbf{x}}|\cdot)$ of synthetic data $\hat{\mathbf{x}}$, similar to the input data distribution $p({\mathbf{x}})$.
Conditional generation further utilizes the paired condition ${\mathbf{y}}$ to guide time series generation of each data sample ${\mathbf{x}}$ as $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. Thus, we try to make the conditional distribution $p(\hat{\mathbf{x}}|{\mathbf{y}})$ of synthetic data $\hat{\mathbf{x}}$ closely mirror the conditional distribution of input data $p({\mathbf{x}}|{\mathbf{y}})$. Considering that the condition in time series is often also a time series as ${\mathbf{y}}\in\mathcal{X}$, Coletta et al. (2024) divided the constraints into two types: (i) soft constraint to guide the time series generation such as trend, or (ii) hard constraint where we should follow during synthesis such as fixed points. In a broader view, well-known time series tasks can be considered as conditional generation, e.g., the unmasked part in imputation or the historical data of forecasting as conditions.
With the diffusion bridge framework, we hypothesize that the standard Gaussian prior might not be the optimal choice. Thus, we investigate a general time series synthesis framework by analyzing the prior selection in diffusion models based on the following research questions (RQs):

### v2

# TimeBridge: Better Diffusion Prior Design with Bridge Models for Time Series Generation

@@ Abstract.
Time series generation is widely used in real-world applications such as simulation, data augmentation, and hypothesis testing. Recently, diffusion models have emerged as the de facto approach to time series generation, enabling diverse synthesis scenarios. However, the fixed standard-Gaussian diffusion prior may be ill-suited for general time series data, such as temporal order and fixed points. In this paper, we propose TimeBridge, a framework that flexibly synthesizes time series data by using diffusion bridges to learn paths between a chosen prior and the data distribution. We then explore several prior designs tailored to time series synthesis. Our framework covers (i) data- and time-dependent priors for unconditional generation and (ii) scale-preserving priors for conditional generation. Experiments show that our framework with data-driven priors outperforms standard diffusion models on time series generation.
Time Series Generation, Diffusion Bridge, Diffusion Prior
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06conference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYccs: Computing methodologies Modeling methodologiesccs: Mathematics of computing Time series analysis
1. Introduction
Synthetic time series have expanded their applications by serving roles in data augmentation (Wen et al., 2021), data privacy (Wang et al., 2024), data compression (Zhou et al., 2023), and hypothesis‐driven simulations of data streams (Xia et al., 2024). Common tasks such as imputation and forecasting can be treated as special cases of time series generation. With advances in deep generative modeling, practitioners have developed VAE‐based (Naiman et al., 2024), GAN‐based (Yoon et al., 2019), and diffusion‐based methods (Ho et al., 2020; Karras et al., 2022). Diffusion approaches are especially effective, yielding high‐fidelity and diverse samples (Tashiro et al., 2021; Alcaraz and Strodthoff, 2023) by estimating the reverse trajectory of a forward noising process toward a standard Gaussian prior (Song et al., 2021).
@@ 1. Introduction
Nevertheless, diffusion models may fail to capture characteristics of time series data because they rely on a fixed mapping from standard Gaussian noise to the data domain. Such restricted priors do not inherently reflect sequential order or scale (Kim et al., 2022), both of which are crucial for modeling historical statistics and temporal dependencies (Zeng et al., 2023). Adding domain‐specific constraints also remains difficult (Harvey et al., 2022; Du et al., 2023); for instance, Coletta et al. (2024) highlighted that generating data under specific bull or bear market patterns requires tailored embedding or penalty functions for sampling guidance.
@@
Figure 1. Illustration of the proposed time series diffusion bridge framework enabling various prior selections.
Model Sampler Steps Uncond. Gen. Constrained Gen. Decomp. Layer Bridge Flexible Prior
[…]
Recent studies have revisited the noising process in diffusion models, exploring data-dependent forward and reverse distributions (Han et al., 2022; Yu et al., 2024) and joint consideration of diffusion steps with actual timestamps (Biloš et al., 2023; Chen et al., 2023a). We extend this direction through Schrödinger bridge (Schrödinger, 1932) and diffusion bridges (De Bortoli et al., 2021; Zhou et al., 2024), which learn optimal transport paths between the data distribution and a desired prior. By selecting appropriate fixed priors, diffusion bridges have demonstrated strong performance in conditional applications such as image-to-image translation (Zhou et al., 2024) and text-to-speech (Chen et al., 2023b).
In this paper, we propose TimeBridge, a flexible time series generation framework that addresses multiple time series synthesis with the prior design of diffusion bridge, as shown in Figure 1. Then, we ask what priors are best suited to time series generation as a replacement for the standard Gaussian prior, enabling more efficient and higher-quality synthesis across diverse settings. Against standard diffusion baselines compared in Table 1, the proposed TimeBridge outperforms in both unconditional and conditional tasks, achieving better results across several datasets in Figure 2.
Our key contributions with flexible prior selection are:
  • We explore diverse priors to alter the standard Gaussian for capturing time series characteristics and propose a new diffusion bridge model tailored to time series generation.
  • For unconditional generation, we design data- and time-dependent priors that reflect intrinsic temporal properties.
@@
  • For conditional generation, we embed scale constraints into a pairwise prior and propose point-preserving sampling, extending its application to imputation tasks.
@@
Figure 2. Performance measures (lower is better) on different tasks and sampling time (lower is better) of time series diffusion models. The proposed TimeBridge achieves both generalization in various tasks and sampling efficiency.
2. Preliminaries
@@
Consider the multivariate time series ${\mathbf{x}}$ defined within the sample space $\mathcal{X}=\mathbb{R}^{d\times\tau}$, where $d$ represents the dimensionality of the variables and $\tau$ denotes the time length. For each index $i\in\{1,\ldots,N\}$ among the $N$ samples, each time series is denoted as ${\mathbf{x}}^{i}=({\bm{x}}^{i}(1),\ldots,{\bm{x}}^{i}(\tau))\in\mathcal{X}$, given the data ${\bm{x}}^{i}(k)\in\mathbb{R}^{d}$ for each timestamp $k\in\{1,\ldots,\tau\}$. For diffusion steps $t\in\{0,\ldots,T\}$, we denote intermediate time series samples as ${\mathbf{x}}_{t}$, and for a specific instance $i$ as ${\mathbf{x}}_{t}^{i}$.
2.1. Optimal Transport and Schrödinger Bridge
The objective of the Schrödinger Bridge (SB) problem is to find an entropy-regularized optimal transport path between the data distribution $p_{0}$ and a target prior $p_{T}$ by minimizing the KL divergence between paths (Léonard, 2012; Chen et al., 2021):
(1) $\underset{p\in\mathcal{P}_{[0,T]}}{min}D_{KL}(p\lVert p^{ref}),\quad s.t.\ p_{0}=p_{data},\ p_{T}=p_{prior},$
where, $\mathcal{P}_{[0,T]}$ denotes the space of path measures on the interval $[0,T]$, $p^{\text{ref}}$ is a reference path measure, and $p_{0},p_{T}$ are the marginals of $p$. By choosing a forward stochastic differential equation (SDE) associated with the reference process $p^{\text{ref}}$, the SB can be reformulated as following forward-backward SDEs:
(2) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g^{2}(t)\nabla\log\bm{\Psi}_{t}({\mathbf{x}}_{t})]dt+g(t)d{\mathbf{w}}_{t},\quad{\mathbf{x}}_{0}\sim p_{data},\end{split}$
(3) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g^{2}(t)\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},\quad{\mathbf{x}}_{T}\sim p_{prior},\end{split}$
where $\nabla\log\bm{\Psi}_{t}(x_{t})$ and $\nabla\log\hat{\bm{\Psi}}_{t}({\mathbf{x}}_{t})$ are corresponding partial differential equations, detailed in Appendix A.
The Schrödinger bridge problem can be addressed using iterative methods (Kullback, 1968; Fortet, 1940) to simulate trajectories converging to the Schrödinger bridge problem (Vargas et al., 2021; Shi et al., 2024; De Bortoli et al., 2021). However, these iterative methods rely on expensive iterative approximation techniques and have seen limited empirical application (Zhou et al., 2024).
2.2. Diffusion Bridge
Standard Diffusion Models
@@ 2.2. Diffusion Bridge
In standard diffusion models (Ho et al., 2020), the diffusion process is constructed by gradually injecting noise into samples ${\mathbf{x}}_{0}$ drawn from the data distribution $p_{0}$, forwarding them into a fixed standard Gaussian distribution $p_{T}=\mathcal{N}(\bm{0},\bm{I})$. The corresponding forward SDE is:
(4) $d{\mathbf{x}}_{t}={\mathbf{f}}({\mathbf{x}}_{t},t)dt+g(t)d{\mathbf{w}}_{t},$
@@
where ${\mathbf{f}}:\mathbb{R}^{d}\times[0,T]\rightarrow\mathbb{R}^{d}$ represents the drift function, $g:[0,T]\rightarrow\mathbb{R}$ denotes the diffusion coefficient, and ${\mathbf{w}}_{t}$ is the Wiener process. As diffusion models learn to reverse the Equation (4), the reverse SDE (Anderson, 1982) is formulated as follows:
(5) $d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})]dt+g(t)d\bar{{\mathbf{w}}}_{t},$
@@
where $p({\mathbf{x}}_{t})$ refers to the probability density of ${\mathbf{x}}_{t}$. Diffusion models learn to match the score function $\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{t})$ for denoising (Song et al., 2021).
Diffusion Bridge
@@
The aforementioned standard diffusion models cannot be extended to general prior distributions. Instead, we can construct the diffusion process towards any endpoint by applying Doob’s h-transform (Doob and Doob, 1984) to Equation (4). With a fixed endpoint ${\mathbf{y}}$, the transformed process for the diffusion bridge is as
(6) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)+g(t)^{2}{\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)]dt+g(t)d{\mathbf{w}}_{t},\\ {\mathbf{x}}_{0}\sim q_{data}({\mathbf{x}}),{\mathbf{x}}_{T}={\mathbf{y}},\end{split}$
@@
where ${\mathbf{h}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}p({\mathbf{x}}_{T}|{\mathbf{x}}_{t})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}$. For a given data distribution $q_{data}({\mathbf{x}},{\mathbf{y}})$, the diffusion bridge (Schrödinger, 1932; De Bortoli et al., 2021) aims to transport ${\mathbf{x}}_{0}$ to ${\mathbf{x}}_{T}$ where $({\mathbf{x}}_{0},{\mathbf{x}}_{T})=({\mathbf{x}},{\mathbf{y}})\sim q_{data}({\mathbf{x}},{\mathbf{y}})$ while following above forward process. Similar to the standard diffusion model, we can construct the reverse SDE as follows:
(7) $\begin{split}d{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}({\mathbf{s}}({\mathbf{x}}_{t},t,{\mathbf{y}},T)-\\ {\mathbf{h}}({\mathbf{x}}_{t},t,{\mathbf{y}},T))]dt+g(t)d\bar{{\mathbf{w}}}_{t},\end{split}$
@@
(8) ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)=\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})|_{{\mathbf{x}}_{t}={\mathbf{x}},{\mathbf{x}}_{T}={\mathbf{y}}}.$
For training, diffusion bridge learns the score function ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$.
Diffusion bridge models have been investigated due to their ability to flexibly map in various domains, including image-to-image translation (Li et al., 2023; Zhou et al., 2024) and text-to-speech synthesis (Chen et al., 2023b), compared to standard diffusion models. Recently, adapting the advances of standard diffusion, such as consistency (He et al., 2024), and deterministic samplers (Zheng et al., 2025), has been further investigated in the diffusion bridge.
Note that although SB theory covers the concepts of diffusion bridges, we separate SB methods with iterative fitting and score-based diffusion bridge methods for clarity in this paper.
3. Framework
3.1. Scenarios
For time series generation, we cover both unconditional and conditional generations. Unconditional generation considers the total dataset containing $N$ samples, denoted as $\mathcal{D}=\{{{\mathbf{x}}^{i}}\}_{i=1}^{N}$. Our goal is to train a deep generative model to mimic the input dataset and yield synthetic samples statistically similar to a given time series dataset. We aim to make the distribution $p(\hat{\mathbf{x}}|\cdot)$ of synthetic data $\hat{\mathbf{x}}$, similar to the input data distribution $p({\mathbf{x}})$.
Conditional generation further utilizes the paired condition ${\mathbf{y}}$ to guide time series generation of each data sample ${\mathbf{x}}$ as $\mathcal{D}=\{({\mathbf{x}}^{i},{\mathbf{y}}^{i})\}_{i=1}^{N}$. Thus, we try to make the conditional distribution $p(\hat{\mathbf{x}}|{\mathbf{y}})$ of synthetic data $\hat{\mathbf{x}}$ closely mirror the conditional distribution of input data $p({\mathbf{x}}|{\mathbf{y}})$. Considering that the condition in time series is often also a time series as ${\mathbf{y}}\in\mathcal{X}$, Coletta et al. (2024) divided the constraints into two types: (i) soft constraint to guide the time series generation such as trend, or (ii) hard constraint where we should follow during synthesis such as fixed points. In a broader view, the unmasked part in imputation or the historical data in forecasting can be considered as conditions for generation.
3.2. Diffusion Bridge for Time Series
To leverage the prior selections, we propose TimeBridge, a diffusion bridge tailored to time‐series synthesis. Our approach produces high-quality samples with fewer sampling steps and allows flexible prior selection. We now provide a detailed description of the sampler designs, network architecture, and overall framework.
Diffusion Bridge Sampler
@@ 3.2. Diffusion Bridge for Time Series
To approximate the score ${\mathbf{s}}({\mathbf{x}},t,{\mathbf{y}},T)$ in Equation (8), we adopt the Denoising Diffusion Bridge Model (DDBM) sampler (Zhou et al., 2024). Unlike prior time‐series methods based on the Schrödinger bridge (Chen et al., 2023b; Garg et al., 2024), DDBM aligns with recent diffusion frameworks and enables efficient bridge sampling. Accordingly, we exploit the tractable marginal distribution of ${\mathbf{x}}_{t}$ by setting ${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\bm{\varepsilon}}$, where $\alpha_{t}$ and $\sigma_{t}$ are noise‐schedule functions and ${\bm{\varepsilon}}\sim\mathcal{N}(\bm{0},\bm{I})$. Under the variance‐preserving (VP) schedule, the marginal distribution at time $t$ becomes
(9) $q({\mathbf{x}}_{t}|{\mathbf{x}}_{0},{\mathbf{x}}_{T})=\mathcal{N}\left(\hat{\bm{\mu}}_{t},\hat{\sigma}^{2}_{t}\mathbf{I}\right),$
(10) $\hat{\bm{\mu}}_{t}=\textstyle\frac{SNR_{T}}{SNR_{t}}\frac{\alpha_{t}}{\alpha_{T}}{\mathbf{x}}_{T}+\alpha_{t}{\mathbf{x}}_{0}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$
(11) $\hat{\sigma}^{2}_{t}=\textstyle\sigma^{2}_{t}\left(1-\frac{SNR_{T}}{SNR_{t}}\right),$
@@
where signal-to-ratio (SNR) is defined as $SNR_{t}={\alpha^{2}_{t}}/{\sigma^{2}_{t}}$. Since the mean $\hat{\bm{\mu}}_{t}$ is a linear combination of ${\mathbf{x}}_{0}$ and ${\mathbf{x}}_{T}$ at time $t$, the data scale is preserved for translation.
For score matching, Karras et al. (2022) proved that if a denoiser function $D({\mathbf{x}};\sigma)$ minimizes the expected error for samples drawn from $p_{data}$, then $D({\mathbf{x}};\sigma)$ becomes the linear function of score function, i.e., if $D({\mathbf{x}};\sigma)$ minimizes $\mathbb{E}_{{\mathbf{y}}\sim p_{data}}\mathbb{E}_{\bm{{\bm{\varepsilon}}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\bm{I})}\lVert D({\mathbf{y}}+\bm{{\bm{\varepsilon}}};\sigma)-{\mathbf{y}}\rVert^{2}_{2}$, then $\nabla_{\mathbf{x}}\text{log}p(x;\sigma)=(D({\mathbf{x}};\sigma)-{\mathbf{x}})/\sigma^{2}$. Thus, we match $D_{\theta}$ directly, which is equivalent to training the score function (Karras et al., 2022) as
(12) $\nabla_{{\mathbf{x}}_{t}}\text{log}q({\mathbf{x}}_{t}|{\mathbf{x}}_{T})\approx{\mathbf{s}}({D_{\theta}},{\mathbf{x}}_{t},t,{\mathbf{x}}_{T},T)$
@@
Architecture Design
To obtain $D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})$, we follow the backbone of Diffusion-TS (Yuan and Qiao, 2024), which employs an encoder–decoder transformer. Let $w_{(\cdot)}^{i,t}$ denote the input of the interpretable layers for index $i\in{1,\ldots,K}$ at diffusion step $t$. Each decoder block is split into trend and seasonal synthesis layers. First, the trend component is synthesized with a polynomial regressor:
(13) $V_{tr}^{t}=\Sigma_{i=1}^{K}(C\cdot\text{Linear}(w_{tr}^{i,t})+X_{tr}^{i,t}),$
where $X_{tr}^{i,t}$ is the mean value of the output of the $i^{th}$ decoder block with the tensor multiplication ()($\cdot$)( ⋅ ) to $C$.
For seasonal components, we use the Fourier transformation as
(14) $S_{i,t}=\Sigma_{k=1}^{K}A_{i,t}^{\kappa_{i,t}^{(k)}}[cos(2\pi f_{\kappa_{i,t}^{(k)}}\tau c+\Phi_{i,t}^{\kappa_{i,t}^{(k)}})+cos(2\pi\bar{f}_{\kappa_{i,t}^{(k)}}\tau c+\bar{\Phi}_{i,t}^{\kappa_{i,t}^{(k)}})],$
where argTopK is to get the top $K$ amplitudes and $f_{k}$ represents the Fourier frequency of the index $k$ and $\bar{(\cdot)}$ denotes ()($\cdot$)( ⋅ ) of the conjugates. $A_{i,t}^{(k)}$ and $\Phi_{i,t}^{(k)}$ are the phase and amplitude of the $k$-th frequency, respectively, after the Fourier transform $\mathcal{F}$ as
(15) $A_{i,t}^{(k)}=\lvert\mathcal{F}(w_{seas}^{i,t})_{k}\rvert,\Phi_{i,t}^{(k)}=\phi(\mathcal{F}(w_{seas}^{i,t})_{k}),$
(16) $\kappa_{i,t}^{(1)},...,\kappa_{i,t}^{(K)}=\underset{k\in\{1,...,\lfloor\tau/2\rfloor+1\}}{\text{argTopK}}\{A_{i,t}^{(k)}\}.$
Overall Framework
Based on the previous subsections, we can design the output of the denoiser $D_{\theta}$ with both conditioning on noisy data ${\mathbf{x}}_{t}$ dependent on data ${\mathbf{x}}_{0}$ and prior ${\mathbf{x}}_{T}$ as follows:
(17) $D_{\theta}=V_{t}^{tr}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+\sum_{i=1}^{K}S_{i,t}(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T})+R(\theta,{\mathbf{x}}_{t},t,{\mathbf{x}}_{T}),$
where $V_{t}^{tr}$ is the output for the trend synthesis layer, $S_{i,t}$ is the output for each seasonal synthesis layer $i$ among $K$ layers and $R$ is the output for estimated residual. Note that calculating Equation (17) is unique in diffusion bridge since the standard diffusion models (Yuan and Qiao, 2024) do not consider the dependency on prior ${\mathbf{x}}_{T}$ for calculating ${\mathbf{x}}_{t}$.
For training, our objective function $\mathcal{L}_{\theta}$ is formulated as follows:
(18) $\mathcal{L}_{\theta}=\mathbb{E}_{t,{\mathbf{x}}_{0}}[w_{t}(\lVert{\mathbf{x}}_{0}-D_{\theta}({\mathbf{x}}_{t},t,{\mathbf{x}}_{T})\rVert^{2}$
$w_{t}$ indicates the weight scheduler of the loss function, and $\lambda$ indicates the strength of the Fourier transform.
As time series data possess stochasticity (Shen and Kwok, 2023), we directly predict $D_{\theta}$ rather than using the reparametrization technique from noise prediction used in other domains (Karras et al., 2022; Zhou et al., 2024). Moreover, we utilize the second-order Heun sampler (Ascher and Petzold, 1998; Karras et al., 2022), achieving both quality and efficiency over first-order methods.
Overall, we name the proposed framework as TimeBridge, a flexible time series generation method that enables the selection of data-driven priors tailored to time series data based on (i) a flexible diffusion framework with diffusion bridge samplers and (ii) an interpretable decomposition-based time series structure.
4. Prior Design for Time Series Synthesis
Based on the flexibility of choosing priors with the proposed TimeBridge, we hypothesize that the standard Gaussian prior might not be the optimal choice. Thus, we investigate a general time series synthesis framework by analyzing the prior selection in diffusion models based on the following research questions (RQs):